OpenStudy (anonymous):

Find the specified vector or scalar. Show your work. u = -4i + 1j and v = 4i + 1j; Find ‖u+v‖.

OpenStudy (anonymous):

@Directrix

OpenStudy (jango_in_dtown):

u+v=(-4+4)i+(1+1)j=0i+2j so \[||u+v||=\sqrt{0^2+2^2}=2\]

OpenStudy (anonymous):

how do you know if its a vector or scalar?

OpenStudy (jango_in_dtown):

in u and v you have i. j components so its definitely vector but the norm is a scalar quantity.. It determines the scalar length only

OpenStudy (anonymous):

thank you can you help with some more?

OpenStudy (jango_in_dtown):

give me the question. i have 5 mins in my hand

OpenStudy (anonymous):

Find the reference angle for the given angle. Show your work. -404°

OpenStudy (anonymous):

or Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work.

OpenStudy (jango_in_dtown):

you have to solve for x

OpenStudy (jango_in_dtown):

solve and write the answer

OpenStudy (anonymous):

okay what about for the first one?

OpenStudy (jango_in_dtown):

for the first one add 360 to the given angle until it lies in between 0 and 360 degree

OpenStudy (jango_in_dtown):

we have -404+360=-44 -44 doesnot lie in between 0 and 360 so add 360 again -44+360=316 degree

OpenStudy (jango_in_dtown):

@heather8

OpenStudy (jango_in_dtown):

i got to go. bye

OpenStudy (anonymous):

thank you!!